Contributed Talks

• 10:45 Singular systems, state feedback problem Dodig, Marija, CELC, Universidade de Lisboa, Lisbon, Portugal

• 11:10 On Schur complements of H-matrices (with R. Bru, I. Giménez, J. Mas) Corral, Cristina, Universidad Politécnica de Valencia, Valencia, Spain

• 11:35 Commutativity preserving maps on real matrices Fosner, Ajda, Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

• 12:00 A Canonical Form for Quasi-Real Normal Matrices Under Real Orthogonal Similarity (with Geoffrey R. Goodson) Horn, Roger, University of Utah, Salt Lake City, USA

• 12:25 On Generalized Procrustes Problem (with João Xavier) Stosic, Marko, Instituto de Sistemas e Robotica, IST, Lisbon, Portugal

Contributed Talks

• 10:45 Acyclic Birkhoff Polytope (with C.M. da Fonseca and Enide Andrade Martins) Costa, Liliana, Centre for Research on Optimization and Control, Aveiro, Portugal

• 11:10 Rank partitions and covering numbers under small perturbations of an element Fernandes, Rosário, Centro de Estruturas Lineares e Combinatórias (CELC), Lisboa, Portugal

• 11:35 Spectrally Arbitrary Ray Patterns Stuart, Jeffrey, Pacific Lutheran University, Tacoma, USA

• 12:00 Nonsingularity of Divisor Tournaments (with Rohan Hemasinha and Jeffrey L. Stuart) Weaver, James, University of West Florida, Pensacola, FL 32514, USA

• 12:25 Additivity obstructions for integral matrices and pyramids (with Miguel Santoyo) Vallejo, Ernesto, Instituto de Matemáticas, Morelia, México • 16:55 GPEA’s researches about the meta resources in teaching and

learning the notion of basis of a vector space (with Bianchini, B.L. and Maranhão, M.C.S.A.) Machado, Silvia, Pontificia Universidade Catolica de São Paulo, São Paulo, Brasil

• 17:20 Whatever Happened to Rook’s Pivoting?

Poole, George, East Tennessee State University, Johnson City, USA

• 17:45 Students difficulties with concept of vector space from point of view of APOS Theory Vargas, Xaab Nop, ICYTDF, México, México

• 18:10 Generalized inverses of infinite matrices Kopparty, Bhaskara Rao, Indiana State University, Terre Haute, IN, USA

• 18:35 Euclidean Jordan algebras and inequalities on the parameters and on the spectra of a strongly regular graph Vieira, Luis, Feup, Porto, Portugal

Contributed Talks

• 16:55 On a singular Toeplitz pencil Gouveia, María, Department of Mathematics, FCTUC, Coimbra, Portugal

• 17:20 General solution of certain matrix equations arising in filter design applications Domínguez, María Elena, Universidad Politécnica de Madrid, Madrid, Spain

• 17:45 The ratio between the Toeplitz and the unstructured condition number (with H. Sekigawa) Rump, Siegfried M., Hamburg University of Technology, Hamburg, Germany

• 18:10 On some classes of H-matrices (with Ljiljana Cvetković, Vladimir Kostić and Francisco Pedroche) Bru, Rafael, Univ. Politécnica, Valencia, Spain • 18:35 Uniform boundedness of Toeplitz Matrices with variable coef-

ficients Grudsky, Sergey, CINVESTAV del I.P.N., México, México

Contributed Talks

• 16:55 On trace inequalities for products of matrices Furuichi, Shigeru, Nihon University, Tokyo, Japan

• 17:20 About the logarithm function over the matrices Bourgeois, Gerald, Faculté de Luminy, Marseille, France

• 17:45 Inequalities on an indefinite inner product space (with N. Bebiano et al.) Soares, Graça, University of Tras-os-Montes and Alto Douro, Vila Real, Portugal

• 18:10 A New Majorization between functions Uchiyama, Mitsuru, Shimane University, Matsue, Shimane, Japan

• 18:35 Contractions and nonlinear matrix equations on positive definite cones (with Yongdo Lim) Lee, Hosoo, Kyungpook National University, Daegu, Korea

Contributed Talks

• 10:35 Sign Patterns That Require Almost Unique Rank (with Frank Hall, Zhongshan Li, Assefa Merid, Yubin Gao) Arav, Marina, Georgia State University, Atlanta, GA, USA

• 11:00 Complete Orthogonal Decomposition Compared with Direct Projection Methods (with Beatriz Lacruz) Pruneda, Rosa E., University of Castilla-La Mancha, Ciudad Real, Spain

• 11:25 Locally linearly dependent operators Semrl, Peter, University of Ljubljana, Ljubljana, Slovenia

• 11:50 On low rank perturbations of matrices (with Luis Manuel Rivera) Glebsky, Lev, Universidad Autónoma de San Luis Potosi, San Luis Potosi, México • 12:15 The class of Inverse-Positive matrices with checkerboard pat-

tern (with Manuel F. Abad, and Juan R. Torregrosa) Gassó, Maria T., Inst.Mat.Mult. Universidad Politécnica Valencia, Valencia, Spain

Contributed Talks

• 10:35 Controllability of Matrices with Prescribed Blocks Cravo, Glória, University of Madeira and CELC, Funchal, Portugal

• 11:00 Perturbations preserving conditioned invariant subspaces (w ith A. Compta and J. Ferrer) Peña, Marta, Universitat Politecnica de Catalunya, Barcelona, Spain

• 11:25 Pencils with Prescribed Constant Subpencils (with F.C. Silva) Roca, Alicia, Dpto. de Matemática Aplicada, Universidad Politécnica, Valencia, Spain

• 11:50 Reachability of regular switched linear systems Carriegos, Miguel, Universidad de León, León, Spain

• 12:15 Geometric structure of the equivalence classes of a controllable pair (with A. Compta, and M. Peña) Ferrer, Josep, Universitat Politecnica de Catalunya, Barcelona, Spain

Contributed Talks

• 16:55 On some general determinantal identities of Sylvester type (with Michela Redivo-Zaglia) Russo, Maria Rosaria, Department of Mathematics - University of Padua, Padova, Italy

• 17:20 Algorithms for computing the Geronimus Transformation Bueno Cachadina, María Isabel, The University of California at Santa Barbara, Santa Barbara, USA

• 17:45 An iterative method for transport equations in radiotherapy (with Bruno Dubroca) Frank, Martin, University of Kaiserslautern, Kaiserslautern, Germany

• 18:10 A structure-preserving Arnoldi-like method for a class of structured matrices Salam, Ahmed, Université du Littoral-Côte d’Opale, Calais, France • 18:35 Polynomial regression in the Bernstein basis (with Ana Marco)

Martínez, José-Javier, Universidad de Alcalá, Alcalá de Henares, Spain

Contributed Talks

• 10:35 Convex and Non-convex Optimization Problems for the Field of Values of a Matrix Uhlig, Frank, Mathematics, Auburn University, Auburn, AL 36849, USA

• 11:00 Total least-squares regularization of Tykhonov type and an ancient racetrack in Corinth (with Kyle Snow) Schaffrin, Burkhard, Ohio State University, Columbus, OH, USA

• 11:25 Numerical ranges of nilpotent operators (with Hwa-Long Gau) Wu, Pei Yuan, National Chiao Tung University, Hsinchu, Taiwan

• 11:50 p-radii of linear operators and equations of self-similarity Protasov, Vladimir, Moscow State University, Moscow, Russia

• 12:15 Matrix Optimization in Statistics Seddighin, Morteza, Indiana University East, Richmod, Indiana

Contributed Talks

• 10:35 The Gauss-Lucas theorem and the numerical range Wrobel, Iwona, Warsaw University of Technology and Polish Academy of Sciences, Warsaw, Poland

• 11:00 Manifold of proper elements (with S. Djordevic)

Sánchez Perales, Salvador, Benemérita Universidad Autónoma de Puebla, Puebla, México

• 11:25 Order Invariant Spectral Properties for Several Matrices (with Charles Johnson) Furtado, Susana, Centro de Estruturas Lineares e Combinatórias da U. L, Porto, Portugal

• 11:50 Spectral Manifolds (with A. Daniilidis, J. Malick, and A. Lewis) Sendov, Hristo, The University of Western Ontario, London, Canada • 12:15 An inequality for the multiplicity of an eigenvalue

Fonseca, Carlos, Department of Mathematics, University of Coimbra, Coimbra, Portugal

Contributed Talks

• 10:35 On varieties of commuting triples Sivic, Klemen, Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

• 11:00 Homomorphisms of matrix semigroups over division rings from dimension two to three (with Gregor Dolinar) Marovt, Janko, Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

• 11:25 Can One Invert a Matrix via Graph Manipulations? (with Tom Bella and Pavel Zhlobich) Olshevsky, Vadim, University of Connecticut, Storrs, USA

• 11:50 Solution of the linear time-dependent potential by using a solvable Lie algebra^{∗} (with M. Villa, and L. Sandoval) Palma, Alejandro, Instituto de Física (BUAP), Puebla, México

• 12:15 The algebraic relations of curvatures of PL manifolds Satô, Kenzi, Tamagawa University, Tokyo, Japan

Contributed Talks

• 10:35 Using Linear Algebra in Teaching Hamilton Quaternions and Graphs Kortesi, Peter, University of Miskolc, Miskolc, Hungary

• 11:00 Representations of trilinear products in Comtrans algebras (with Jonathan D. H. Smith) Im, Bokhee, Chonnam National University, Gwangju, Korea(Rep. of)

• 11:25 Matrix Results on Weighted Drazin Inverse and Some Applications (with Adem Kilicman) Al Zhour, Zeyad, Zarqa Private University, Zarqa, Jordan

• 11:50 Algorithm resolving problem of determination of finite cyclicgroup by its automorphism group (with V.K. Vildanov and A.L. Sylla) Sebeldin, Anatoly, University UGANC, Guinea, Conakry, Guinea • 12:15 Z_{2}-graded symmetry classes of tensors

Shahryari, Mohammad, Tabriz University, Tabriz, Iran, Tabriz, Iran

Contributed Talks

• 16:55 Graph Energy Change Due to Edge Deletion Day, Jane, San Jose State University, San Jose, CA, USA

• 17:20 Minimum Rank of Graphs with Loops Mikkelson, Rana, Iowa State University, Ames, IA, United States of America

• 17:45 On the equivalence class graph (with Rosário Fernandes) Perdigão, Cecília, Faculdade de Cincias e Tecnologia-UNL, Lisboa, Portugal

• 18:10 Orthogonal Removal of Vertices and Minimum Semidefinite Rank (with Sivaram Narayan) Mitchell, Lon, Virginia Commonwealth University, Richmond, United States

• 18:35 An upper bound for the maximum nullity of a symmetric matrix whose graph is outerplanar Sinkovic, John, Technische Universiteit Eindhoven, Eindhoven, Netherlands

Contributed Talks

• 16:55 Some conditions for the commutativity of matrix patterns (with C.R. Johnson) Marques, Maria-da-Graça, University of Algarve and CELC, Faro, Portugal

• 17:20 General preservers of quasi-commutativity (with Bojan Kuzma) Dolinar, Gregor, Faculty of Electrical Engineering, Ljubljana, Slovenia

• 17:45 An extended Lie-Trotter formula and its applications (with Sejong Kim and Yongdo Lim) Ahn, Eunkyung, Kyungpook National University, Daegu, Korea

• 18:10 On the matrices that preserve the value of the immanant of the upper triangular matrices (with Rosário Fernandes) da Cruz, Henrique F., U.B.I, Covilhã, Portugal • 18:35 Complexity Analysis of the Primal-Dual Interior Point Method

for Second-order Cone Optimization Problem (with Bo Kyung Choi) Lee, Gue Myung, Pukyong National University, Busan, Korea

Contributed Talks

• 16:55 On the problem of diagonalizability of matrices over a principal ideal domain Prokip, Volodymyr, Institute of Appl.Problem for Mech.and Math. NAS of Ukr, L’viv, Ukraine

• 17:20 Euler’s difference table and maximum permanents of (0, 1)- matrices Rakotondrajao, Fanja, University of Antananarivo, Antananarivo, Madagascar

• 17:45 Shrinkage Estimators for Estimation the Multivariate Normal Mean Vector under Degrees of Distrust (with Abdunnabi M. Ali Elbouzedi) Hanaish, Ibrahim, 7th October University, Misurata, Libya

• 18:10 Asymptotic Behavior of Solutions of Stochastic Equations and Applications in Statistical Parameter Estimation Kaibah, Hussein, 7th October University, Misurata, Libya

Contributed Talks

• 10:35 One upper bound for the largest eigenvalue of the signless Laplacian (with Slobodan Simic) Andjelic, Milica, Center for Research on Optimization and Control, Aveiro, Portugal

• 11:00 On the spectra of some graphs like weighted rooted trees (with Rosário Fernandes and Helena Gomes) Martins, Enide, Centre for Research on Optimization and Control (CEOC), Aveiro, Portugal

• 11:25 Locally computable approximations of absorption times for graph clustering (with Pekka Orponen and Vanesa Avalos) Schaeffer, Elisa, Universidad Autónoma de Nuevo León, San Nicolás de los Garza, México

• 11:50 The Laplacian Spectrum of Hypergraphs (with Aida Kamisalic) Rodríguez, Juan Alberto, Universitat Rovira i Virgili, Tarragona, España • 12:15 Linearly Independent Vertices and Minimum Semidefinite Rank

Narayan, Sivaram, Central Michigan University, Mount Pleasant, Michigan 48859, USA

Contributed Talks

• 10:35 Further results on the representation of the Drazin inverse of a 2 × 2 block matrix (with M.F. Martínez-Serrano and N. Castro-González) Dopazo, Esther, Facultad de Informatica. Universidad Politécnica Madrid, Boadilla del Monte, Madrid, Spain

• 11:00 Duality in the Hopf Algebra of multivariate polynomials Bengochea, Gabriel, Universidad Autónoma de la Ciudad de México, México D.F.

• 11:25 Representations for the generalized Drazin inverse of additive perturbations (with M.F. Martínez-Serrano) Castro-González, Nieves, Universidad Politécnica de Madrid, Madrid, Spain

• 11:50 Some additive results on Drazin Inverses (with R.E. Hartwig) Patricio, Pedro, Departamento de Matemática, Universidade do Minho, Braga, Portugal

• 12:15 Linear algebraic approach to rational functions Verde-Star, Luis, UAM Iztapalapa, México DF, México

Contributed Talks

• 15:30 Some properties of the class sign regular matrices and its subclasses (with J.M. Peña) Cortés, Vanesa, Universidad de Zaragoza, Zaragoza, Spain

• 15:55 A Completion Problem over the Field of Real Numbers (with Fernando C. Silva) Teixeira Matos, Isabel, Centro de Estruturas Lineares e Combinatórias (CELC), Lisboa, Portugal

• 16:20 An upper bound on the characteristic polynomial of a nonnegative matrix leading to a proof of the Boyle–Handleman conjecture (with Michael Neumann) Goldberger, Assaf, Tel Aviv University, Tel Aviv, Israel • 16:45 Orderings of matrix algebras and their applications

Wojciechowski, Piotr, University of Texas at El Paso, El Paso, USA

• 17:10 On The Root Clustering of Matrices Esen, Özlem, Anadolu University, Eskisehir, Turkey

Contributed Talks

• 15:30 Reduction of an Ito’s diffusion input output model for the determination of square mean stability Guzmán, José Ramón, Instituto de Investigaciones Económicas. UNAM., México, México

• 15:55 On Euclidean Squared Distance Matrices (with Chi-Kwong Li and Michael Trossett) Milligan, Thomas, University of Central Oklahoma, Edmond, OK, USA

• 16:20 The strong closure of the similarity orbit for a class of pairs of finite rank operators Barría, José, Santa Clara University, Santa Clara, U.S.A.

• 17:10 One Type of Inverse Eigenvalue Problems in Quaternionic Quantum Mechanics (with Liang-gui Feng) Cheng, Wei, National University of Defense Technology, Changsha, P.R.China